# Stochastic Multi-round Submodular Optimization with Budget

### v1

@@ Abstract
In this work we study the problem of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm), in which we would like to maximize the sum over multiple rounds of the value of a monotone and submodular objective function, subject to the fact that the values of this function depend on the realization of stochastic events and the number of observations that we can make over all rounds is limited by a given budget. This problem extends, and generalizes to multiple round settings, well-studied problems such as (adaptive) influence maximization and stochastic probing.
We first show that whenever a certain single-round optimization problem can be optimally solved in polynomial time, then there is a polynomial time dynamic programming algorithm that returns the same solution as the optimal algorithm, that can adaptively choose both which observations to make and in which round to have them. Unfortunately, this dynamic programming approach cannot be extended to work when the single-round optimization problem cannot be efficiently solved (even if we allow it would be approximated within an arbitrary small constant). Anyway, in this case we are able to provide a simple greedy algorithm for the problem. It guarantees a $(1/2-\epsilon)$-approximation to the optimal value, even if it non-adaptively allocates the budget to rounds.
@@ 1 Introduction
Consider the well-known problem of influence maximization: here we are given a (social) network, with edge probabilities encoding the strength of the relationships among network’s components; these probabilities describe how easily influence (e.g., an information or an innovation) spreads over the network starting from a limited set of nodes, usually termed seeds; the goal is then to find the set of seeds that maximizes the number of “influenced” nodes. The non-adaptive version of the problem, in which all the seeds are chosen before that influence starts, has been widely studied, with numerous successful algorithms [26, 10, 7] and applications ranging from viral marketing [26], to election manipulation [33, 13, 8], to infective disease prevention [34]. Recently, focus moved to the adaptive variant of this problem (in which seeds can be added even when the influence is spreading from previously placed seeds) [22, 28, 14]. This variant happens to have strong relationship with some Stochastic Probing problems [3], in which there are $n$ items whose states are stochastically defined according to (not necessarily equal) independent distributions and the goal is to choose a feasible subset of items maximizing the value of a non-negative, monotone and submodular set function. Several stochastic probing settings can be also encountered in expert hiring [24], online dating and kidney exchange [23], Bayesian mechanism design [2], and sequentially posted pricing [9, 1].
Research both in influence maximization and stochastic probing focused almost exclusively on the single-round setting. Recently, there is an increasing interest in those cases in which the optimization is run in a multi-round setting. This is highlighted both by the increasing amount of literature on the topic (see Related Works), and by the numerous applications that are evidently deployed over multiple rounds, as it is the case, e.g., for many real-word marketing or electoral campaigns, or expert hiring sessions. Moreover, it is often the case that in these marketing, electoral, or hiring campaigns we are interested both in maximizing the total effectiveness of the campaign, and in limiting the total budget spent. Note that, in order to achieve this goal it may be insufficient to limit the budget spent at each round, since this may lead in general to very inefficient outcomes (see Remark 1 below for more details). However, this makes the problem much more complex: we do not only need to care about which action to take at each round, but also we need to care about how large must be the portion of the budget spent at that round.
In this work, we aim to extend the study of above cited problems (and their generalizations) by considering all these properties: multiple-rounds, total budget, and adaptivity.
@@
The interest in multi-round variants of the above discussed problems is testified by the increasing amount of papers focusing on this aspect: for example, Sun et al. [29] considered an influence maximization problem over multiple and independent rounds of diffusion, where the goal is to maximize the number of nodes selected in at least one round, subject to the selection of at most $k$ seeds for each round (thus, in contrast to our setting, they have a limit only on per-round budget); in the work of Tong et al. [30], instead, a round denotes a time step in which either a node is selected, or the already active nodes influence some of their neighbors, and the goal is to select at most $k$ seeds that maximize the influence spread within $T$ rounds.
@@
Moreover, running a campaign over multiple rounds can be very useful (maybe necessary) to the designer of the campaign in order to gain the necessary knowledge about the environment to address the desired goal. This approach has been widely studied under the lens of online learning [21]. Specifically, online learning algorithms for some submodular maximization processes (including the simplest form of our multi-round influence maximization problem) have been proposed [31, 12, 32]. In particular, these works introduced an approach, named Combinatorial Upper Confidence Bound (CUCB), that extended the well-known UCB bandit algorithm [4] with the aim of designing online learning algorithms with good guarantees for a wide class of optimization problems. However, such approach has been defined only for non-adaptive optimization frameworks, e.g., standard influence maximization problems in which the diffusion is not observed after each seed selection, but only at the end of the selection process. Gabillon et al. [20] presented CUCB algorithms for a class of adaptive optimization problems simpler than the one considered in this work, since, e.g., the seeding is subject to a limit on per-round budget and the underlying combinatorial structure is not sufficiently rich to represent influence maximization problems. Budget constraints without any limit on per-round quotas have been considered by Das et al. [16], that still achieves good guarantees, but considers an even simpler setting than the one by Gabillon et al [20], that in fact does not apply to influence maximization.
@@
The analysis of optimization advantages of adaptive w.r.t. non-adaptive algorithms has been initiated by Dean et al [18, 17] on classical packing problems. Only recently, this kind of analysis has been applied to stochastic maximization problems, including single-round influence maximization and stochastic probing. In particular, Golovin and Krause [22] showed that whenever the objective function satisfies a property named adaptive submodularity (see Definition 2 below), a greedy algorithm achieves a $(1-\frac{1}{e})$-approximation of the optimal greedy algorithm. Chen and Peng [28, 11] also proved constant bound on the adaptivity gap for some influence maximization problems, i.e., by using a non-adaptive optimal algorithm we lose only a constant approximation factor w.r.t. the optimal adaptive algorithm. Some bounds on the approximation ratio of adaptive algorithms and on the adaptivity gap have been recently improved by D’Angelo et al [14, 15]. While our model and results are clearly inspired by these works, machineries defined therein cannot be directly applied to our setting since we are also interested in allocating the budget among multiple rounds.
@@
In order to provide results that hold for both influence maximization and stochastic probing (and for variants or generalizations of them), we will first introduce the framework of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm) problems. These problems are defined by a number $T$ of rounds; a ground set of $n$ items, whose state is drawn at each time step from an item-specific set according to some time-dependent distributions; a t; and a budget $B$ on the possible observations done within the time-horizon $T$. The problem is to choose which observations to make in each of the $T$ rounds, subject to the budget limit, in order to maximize the expected total value of the state function over the $T$ rounds.
@@
This problem extends and generalizes both the influence maximization and stochastic probing problems (more details will be provided below): in the first case, the state of an item $v$ defines which nodes would be influenced if $v$ is selected as a seed, with the distribution of the state of $v$ depending on edges’ probabilities; an observation consists on choosing a node as a seed, running the influence spread, and evaluating the realized state by, e.g., counting the number or the total weight of influenced nodes. In the second case, the state of an item can be active or inactive, and we can observe the state of selected items only; the value of the submodular set function depends on the set of selected items which are active. Note that SBMSm not only generalizes influence maximization and stochastic probing problems to multiple round settings, but it also allows to model more realistic settings where states can change over time: e.g., in the influence maximization setting, we may allow edge probabilities or node weights to change in different rounds or be discounted over time.
@@
We will start by characterizing when the problem can be optimally solved. In particular, we will show that whenever a certain single round optimization problem can be optimally solved in polynomial time, then there is a polynomial time dynamic programming algorithm that returns the optimal allocation of observations. Note that this optimal allocation is assumed to be very strong: at each time step, based on observations both in the past and in the current round, it decides if to spend a further unit of budget in the current round to make a new observation, or to stop the observations for the current round, and save budget for next rounds. Note that this will allow the optimum to react immediately to very negative events (e.g., the currently chosen seeds have influenced much less nodes than expected) or to very positive events (e.g., the currently chosen seeds have influenced much more nodes than expected). Hence, we find it is quite surprising that the optimality of the single round problem straightly extends to multiple rounds.
@@
Unfortunately, the single-round problem turns out to usually be a hard problem to solve. Indeed, it can be easily reduced from the (deterministic) monotone submodular maximization problem, that is is known to be NP-hard [19], and no algorithm is known that approximate this problem in general instances within a factor better than $1-1/e$. Nevertheless, we show that there are non-trivial instances for which the problem can be effectively solved in polynomial time, essentially corresponding to the cases in which many observations will provide “sure” results, and essentially few of them are subject to randomness. A natural direction would be to extend our framework so that it would work also with similar but more general instances where, for example, one may compute an arbitrarily good approximation in polynomial time. However, we will show that within our dynamic programming framework this approximation ratio will explode with the number of rounds, and hence fails to provide good approximation guarantees.
For this reason, we need to resort to a different framework. Interestingly, we will show that a greedy polynomial time algorithm, that is arguably simpler than the previous one, provides a $(1/2-\epsilon)$-approximation of the optimal allocation for every instance. Hence, its approximation is very close to what happens in the much simpler single-round case [22].
@@
Interestingly, this greedy approximation algorithm relies on a result that can be of independent interest. Indeed, we prove that in a SBMSs problem, a non-adaptive allocation of budget to rounds can be sufficient to guarantee a constant approximation of the adaptive optimal allocation of observations described above. Even more surprisingly, we can achieve such result by allocating the budget to rounds greedily, i.e., by assigning each unit of budget to the round in which it is expected to produce the largest expected value of the score function, given the amount of budget still available.
@@
Our ($1/2-\epsilon$)-approximate algorithm then follows by composing this greedy allocation of budget, with the algorithm that chooses the items to observe up to the allocated budget in a greedy adaptive way, i.e., it chooses an item that is expected to achieve the maximum value of the state function, observes the realization of the state for this item, and then it chooses the next item among the ones that are expected to maximize the state function given the state realization for the first item, and so on. In other words, our result shows that the powerful optimal allocation of the observations described above, that is adaptive both in the allocation of budget to rounds and in the allocation of observations within rounds, can be well-approximated by an algorithm that is only partially adaptive, i.e., it is adaptive in the allocation of observations within rounds, but not in the allocation of budget to rounds. We point out that the effectiveness of a partially adaptive approach for multi-round adaptive optimization has been left as prominent research direction by [5], and to the best of our knowledge, our results are the first that consider such an approach and guarantee, at the same time, a good approximation to the optimal solution in polynomial time.
@@ 2.1 A Setting of Multi-round Maximization
We consider a setting $I:=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ whose parameters and connected properties are described as follows. $T\geq 2$ is a finite and integral horizon, where each $t\in[T]$ denotes the $t$-th round. $B\geq 0$ is a given budget. $V$ is a ground set of $n\geq 2$ items.
For any fixed $v\in V$, let $H(v)$ be a set of possible local states for item $v$, and let $H\subseteq\times_{v\in V}H(v)$ be the set of possible global states, that is, a global state $\eta\in H$ is a collection of local states $(\eta(v))_{v\in V}$ (with $\eta(v)\in H(v)$ for any $v\in V$). In the following, the global states may be simply called states.
For any round $t$, a state $\eta_{t}\in H$ is picked according to a probability distribution $\mathcal{P}_{t}$ over $H$, that is called (global) state distribution at round $t$; the state distributions $\mathcal{P}_{1},\ldots,\mathcal{P}_{T}$ are not generally independent. In the following, we will adopt bold letters to denote the random variables, and non-bold letters to denote their realization (e.g., $\bm{\eta}(v)$ is a random local-state, and $\eta(v)$ denotes a possible realization of $\bm{\eta}(v)$).
Finally, for any round $t\in[T]$, $f_{t}:2^{V}\times H\rightarrow{\mathbb{R}}_{\geq 0}$ is the state function at round $t$, that associates a non-negative real-value $f_{t}(S_{t},\eta_{t})$ with each subset of items $S_{t}\subseteq V$ and state $\eta_{t}$. Note that state functions are allowed to be different at each time step. Such a variability in both the state distribution and state function, immediately makes inefficient to allocate the same budget to each round: indeed, it is immediate to build a setting in which many rounds may lead to very poor outcomes, and only few of them may lead to a very large outcome, and hence all the budget allocated to the former rounds is wasted.
Let $\bar{f}:(2^{V}\times H)^{T}\rightarrow{\mathbb{R}}_{\geq 0}$ be the aggregated state function, that assigns the value $\bar{f}(\vec{S},\vec{\eta}):=\sum_{t=1}^{T}f_{t}(S_{t},\eta_{t})$ to each aggregated subset $\vec{S}:=(S_{1},\ldots,S_{T})$ and aggregated state $\vec{\eta}:=(\eta_{1},\ldots,\eta_{T})$, i.e., it returns the cumulative value of the state function $f$ over all rounds.
We consider then the problem of selecting an aggregated subset $\vec{S}=(S_{1},S_{2},\ldots,S_{T})$ such that $\bar{f}(\vec{S},\vec{\eta})$ is maximized subject to the budget constraint $\sum_{1}^{T}|S_{t}|\leq B$, for a fixed aggregated state $\vec{\eta}$ drawn from the aggregated state distributions $(\bm{\eta}_{1},\ldots,\bm{\eta}_{T})$. Unfortunately, we need to solve this problem without knowing the aggregated state, and we have access to the aggregated state distributions and functions only.

### v2

@@ Abstract
In this work we study the problem of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm), in which we would like to adaptively maximize the sum over multiple rounds of the value of a monotone and submodular objective function defined on a subset of items, subject to the fact that the values of this function depend on the realization of stochastic events and the number of items that we can select over all rounds is limited by a given budget. This problem extends, and generalizes to multiple round settings, well-studied problems such as (adaptive) influence maximization and stochastic probing.
We first show that, if the number of items and stochastic events is somehow bounded, there is a polynomial time dynamic programming algorithm for SBMSm. Then, we provide a simple greedy approximation algorithm for SBMSm, that first non-adaptively allocates the budget to be spent at each round, and then greedily and adaptively maximizes the objective function by using the budget assigned at each round. Such algorithm guarantees a $(1-1/e-\epsilon)$-approximation to the optimal adaptive value.
Finally, by introducing a metric called budget-adaptivity gap, we measure how much an optimal policy for SBMSm, that is adaptive in both the budget allocation and item selection, is better than an optimal partially adaptive policy that, as in our greedy algorithm, determined the budget allocation in advance. We show a tight bound of $e/(e-1)$ on the budget-adaptivity gap, and this result implies that our greedy algorithm guarantees the best approximation among all partially adaptive policies.
Keywords: Combinatorial Optimization; Stochastic Adaptive Optimization; Submodularity.
@@ 1 Introduction
Consider the well-known problem of influence maximization: here we are given a (social) network, with edge probabilities encoding the strength of the relationships among network’s components; these probabilities describe how easily influence (e.g., an information or an innovation) spreads over the network starting from a limited set of nodes, usually termed seeds; the goal is then to find the set of seeds that maximizes the expected number of “influenced” nodes. The non-adaptive version of the problem, in which all the seeds are chosen before that influence starts, has been widely studied, with numerous successful algorithms [32, 12, 8] and applications ranging from viral marketing [32], to election manipulation [39, 16, 10], campaigns to contrast misinformation [2] to infective disease prevention [40]. Recently, focus moved to the adaptive variant of this problem (in which seeds can be added even when the influence is spreading from previously placed seeds) [26, 34, 17]. This variant happens to have strong relationship with some Stochastic Probing problems [4, 3], in which there are $n$ items whose states are stochastically defined according to (not necessarily equal) independent distributions and the goal is to choose a feasible subset of items maximizing the expected value of a non-negative, monotone and submodular set function. Several stochastic probing settings can be also encountered in expert hiring [30], online dating and kidney exchange [27, 9], Bayesian mechanism design and sequentially posted pricing [11, 1].
Research both in influence maximization and stochastic probing focused almost exclusively on the single-round setting. Recently, there is an increasing interest in those cases in which the optimization is run in a multi-round setting. This is highlighted both by the increasing amount of literature on the topic (see Related Works), and by the numerous applications that are evidently deployed over multiple rounds, as it is the case, e.g., for many real-word marketing, electoral and misinformation/disinformation contrasting campaigns. Moreover, it is often the case that in these contexts we are interested both in maximizing the total effectiveness of the campaign, and in limiting the total budget spent. Note that, in order to achieve this goal it may be insufficient to limit the budget spent at each round, since this may lead in general to very inefficient outcomes (see Remark 1 below for more details). However, this makes the problem much more complex: we do not only need to care about which actions to take at each round, but also we need to care about how large must be the portion of the budget spent at that round.
In this work, we aim to extend the study of above cited problems (and their generalizations) by jointly considering all these properties: multiple-rounds, total budget, and adaptivity.
@@
The interest in multi-round variants of the above discussed problems is testified by the increasing amount of papers focusing on this aspect: for example, [35] considered an influence maximization problem over multiple and independent rounds of diffusion, where the goal is to maximize the number of nodes selected in at least one round, subject to the selection of at most $k$ seeds for each round (thus, in contrast to our setting, they have a limit only on per-round budget); in the work of [36], instead, a round denotes a time step in which either a node is selected, or the already active nodes influence some of their neighbors, and the goal is to select at most $k$ seeds that maximize the influence spread within $T$ rounds.
@@
Moreover, running a campaign over multiple rounds can be very useful (maybe necessary) to the designer of the campaign in order to gain the necessary knowledge about the environment to address the desired goal. This approach has been widely studied under the lens of online learning [25]. Specifically, online learning algorithms for some submodular maximization processes (including the simplest form of our multi-round influence maximization problem) have been proposed [37, 15, 38]. In particular, these works introduced an approach, named Combinatorial Upper Confidence Bound (CUCB), that extended the well-known UCB bandit algorithm [5] with the aim of designing online learning algorithms with good guarantees for a wide class of optimization problems. However, such approach has been defined only for non-adaptive optimization frameworks, e.g., standard influence maximization problems in which the diffusion is not observed after each seed selection, but only at the end of the selection process. [24] presented CUCB algorithms for a class of adaptive optimization problems simpler than the one considered in this work, since, e.g., the seeding is subject to a limit on per-round budget and the underlying combinatorial structure is not sufficiently rich to represent influence maximization problems. Budget constraints without any limit on per-round quotas have been considered by [19], that still achieves good guarantees, but considers an even simpler setting than the one by [24], that in fact does not apply to influence maximization.
@@
The analysis of optimization advantages of adaptive w.r.t. non-adaptive algorithms has been initiated by [21, 20] on classical packing problems, mainly trough the adaptivity gap, a metric defined as the worst-case ratio between the value achieved by an optimal adaptive algorithm and the optimal non-adaptive value. Only recently, this kind of analysis has been applied to stochastic maximization problems, including single-round influence maximization and stochastic probing. In particular, [26] showed that whenever the objective function satisfies a property named adaptive submodularity (see Definition 2 below), an approximate greedy algorithm achieves a $(1-1/e-\epsilon)$-approximation of the optimal greedy algorithm, with $e\approx 2.718$ denoting the Nepero number (that is, $1-1/e\approx 0.632$) and $\epsilon>0$ being an arbitrarily small constant. [34, 13] also proved constant bound on the adaptivity gap for some influence maximization problems, i.e., by using a non-adaptive optimal algorithm we lose only a constant approximation factor w.r.t. the optimal adaptive algorithm. Some bounds on the approximation ratio of adaptive algorithms and on the adaptivity gap have been recently improved by [17, 18, 14]. The adaptivity gap of stochastic probing problems has been also extensively studied [4, 3, 28, 29, 9].
While our model and results are clearly inspired by these works, machineries defined therein cannot be directly applied to our setting since we are also interested in allocating the budget among multiple rounds (for further details see, for instance, Remark 2 below).
@@
In order to provide results that hold for both influence maximization and stochastic probing (and for variants or generalizations of them), we will first introduce the framework of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm) problems. These problems are defined by the following quantities: a number $T$ of rounds; a ground set $V$ of $n$ items, whose state is drawn at each time step from an item-specific set according to some time-dependent distributions; a monotone submodular objective function $f_{t}$ at each round $t$, defined on subsets of $V$; a budget $B$ on the possible selectable items within the time-horizon $T$. The problem is to choose which items to select in each of the $T$ rounds, subject to the budget limit, in order to maximize the expected total value of the objective functions over the $T$ rounds. Furthermore, the items can be selected in an adaptive way, that is, after selecting an item, one observes the related state and the new value of function $f_{t}$, and based on this information one can decide whether to select other items in the current round, or whether to save the budget to select items in subsequent rounds.
@@
This problem extends and generalizes both the influence maximization and stochastic probing problems (more details will be provided below). In the first case, an item is a node $v$ that can be selected as a seed, the state of an item $v$ defines which nodes would be influenced if $v$ is a seed, with the distribution of the state of $v$ depending on edges’ probabilities, and the objective function counts the number or the total weight of influenced nodes. In the second case, the state of an item can be active or inactive, and we can observe the state of selected items only; the value of the submodular set function depends on the set of selected items which are active. Note that SBMSm not only generalizes influence maximization and stochastic probing problems to multiple round settings, but it also allows to model more realistic settings where states can change over time: e.g., in the influence maximization setting, we may allow edge probabilities or node weights to change in different rounds or be discounted over time.
@@
We will start by characterizing when the problem can be optimally solved. In particular, we first define a certain optimization problem parametrized by a round $t\leq T$ and a budget $b\leq B$, called Single-Round Maximization ($\text{SRm}(t,b)$). Then we show that, whenever $\text{SRm}(t,b)$ can be optimally solved in polynomial time, there is a polynomial time dynamic programming algorithm that returns the optimal solution for the general multi-round problem SBMSm. Note that the optimal adaptive selection of items over a multiple-round window, as required by SBMSm, is assumed to be very strong: at each time step, based on observations both in the past and in the current round, the algorithm decides if to spend a further unit of budget in the current round to select a new item, or to stop the selection of items for the current round, and save budget for next rounds. Note that this will allow the optimum to react immediately to very negative events (e.g., the currently chosen seeds have influenced much less nodes than expected) or to very positive events (e.g., the currently chosen seeds have influenced much more nodes than expected). Hence, we find it is quite surprising that the optimality of the single-round problem straightly extends to multiple rounds.
Unfortunately, $\text{SRm}(t,b)$ turns out to usually be a hard problem to solve. Indeed, it can be easily reduced from the (deterministic) monotone submodular maximization problem, that is known to be APX-hard [23], and there is no algorithm that approximates this problem in general instances within a factor better than $1-1/e$ (unless P=NP). Nevertheless, we show that there are non-trivial instances for which the problem can be effectively solved in polynomial time, essentially corresponding to the cases in which there are a few items and the possible states of the underlying stochastic environment are polynomial in the size of the input instance. The obtained findings imply that both $\text{SRm}(t,b)$ and SBMSm are fixed-parameter-tractable (FTP) problems [22] w.r.t. to the number of items, if the outcomes of the stochastic environment are polynomially bounded.
@@
A natural direction would be to extend our framework so that it would work also with similar but more general instances where, for example, one may compute an arbitrarily good approximation in polynomial time. However, within our dynamic programming framework, this approximation ratio would explode with the number of rounds, and hence fails to provide good approximation guarantees.
For this reason, we need to resort to a different framework. Interestingly, we will show that a greedy polynomial time algorithm, that is arguably simpler than the previous one, provides a $(1-1/e-\epsilon)$-approximation of the optimal adaptive solution for every instance, where $\epsilon>0$ is an arbitrarily small constant due to the use of some randomized sampling procedures. Hence, the obtained approximation matches that guaranteed in the much simpler single-round case by [26] and cannot be substantially improved in light of the hardness result of [23]; we also point that the machinery of [26] cannot be directly applied to our multi-round problem to obtain the same approximation guaranteed for the restricted single-round setting (see Remark 2 below), and this further made the introduction of a new and more general algorithmic approach necessary.
@@
Interestingly, our greedy approximation algorithm relies on a result that can be of independent interest. Indeed, we prove that in a SBMSs problem, a non-adaptive allocation of budget to rounds can be sufficient to guarantee a constant approximation of the optimal adaptive solution. Even more surprisingly, we can achieve such result by allocating the budget to rounds greedily, i.e., by assigning each unit of budget to the round in which it is expected to produce the largest expected value of the objective function, given the amount of budget still available. Our ($1-1/e-\epsilon)$-approximate algorithm then follows by composing this greedy allocation of budget, with the algorithm that chooses the items to observe up to the allocated budget in a greedy adaptive way, i.e., it chooses an item that is expected to achieve the approximate maximum value of the objective function with high probability (with both approximation and probability depending on the choice of $\epsilon>0$), observes the realization of the state for this item, and then it chooses the next item among the ones that are expected to maximize the objective function given the state realization for the first item, and so on. In other words, our result shows that the powerful optimal adaptive solution for an SBMSs problem, that is adaptive both in the allocation of budget to rounds and in the choice of selected items within rounds, can be well-approximated by an algorithm that is only partially adaptive, i.e., it is adaptive in the choice of selected items, but not in the allocation of budget to rounds.
Apart from the good approximation guarantees achieved by our greedy algorithm, the non-adaptive budget allocation is of independent interest, as it makes sense for several multi-round settings arising, for instance, from marketing or political campaigns. In fact, assuming that the considered campaign must be planned in multiple periods (i.e. rounds), it may be preferable to allocate the budget to be spent for each period in advance, due to organizational and management reasons. Therefore, for this and other settings, it would be good to measures how much an optimal policy for SBMSm, that is adaptive in both the budget allocation and the choice of the selected items, is much better than a simpler partially adaptive policy that, instead, determines the budget to be spent for each round in a non-adaptive way (that is, before starting the adaptive selection of the items and without observing their realizations).
To this aim, we introduce a metric called budget-adaptivity gap (see Definition 3 below), defined as the worst-case ratio between the fully optimal adaptive value and that guaranteed by a partially adaptive policy that non-adaptively allocates the budget. The notion of budget-adaptivity gap can be seen as a novel and interesting variant of the standard adaptivity gap [21, 20], that compares optimal adaptive strategies with fully non-adaptive ones, mainly for single-round optimization problems. Indeed, our variant still estimates the impact of non-adaptivity, but on the budget allocation only, that is a fundamental aspect to be taken into account to adaptively maximize the objective functions in multi-round settings.
@@
We show that the budget-adaptivity gap is exactly $e/(e-1)\approx 1.582$, that is, the non-adaptive budget allocation, in general, cannot be used to reach the fully adaptive optimum but still constitutes a good approximation of the fully adaptive ones; this result also implies that the above greedy algorithm essentially guarantees the best approximation among all partially adaptive policies. We point out that the effectiveness of a partially adaptive approach for multi-round adaptive optimization has been left as prominent research direction by [6], and to the best of our knowledge, our results are the first that consider such an approach and guarantee, at the same time, a good approximation to the optimal solution in polynomial time.
@@ 2.1 A Setting of Multi-round Maximization
We consider a setting $I:=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ whose parameters and connected properties are described as follows.
$T\geq 2$ is a finite and integral horizon, where each $t\in[T]$ denotes the $t$-th round. $B\geq 0$ is a given budget. $V$ is a ground set of $n\geq 2$ items. For any fixed $v\in V$, let $H(v)$ be a set of possible local states for item $v$, and let $H:=\times_{v\in V}H(v)$ be the set of global states, that is, a global state $\eta\in H$ is a collection of local states $(\eta(v))_{v\in V}$ (with $\eta(v)\in H(v)$ for any $v\in V$). In the following, the global states may be simply called states.
For any round $t$, a state $\eta_{t}\in H$ is picked according to a probability distribution $\mathcal{P}_{t}$ over $H$, that is called (global) state distribution at round $t$. The state distributions $\mathcal{P}_{1},\ldots,\mathcal{P}_{T}$ are assumed to be independent; however, for any fixed round $t\in[T]$, the probability distributions induced by local states $\eta_{t}(v)$ from $\mathcal{P}_{t}$ may be dependent. We observe that, even if we assumed that each tuple $\eta:=(\eta(v))$ with $\eta(v)\in H(v)$ is a global state, we can impose the unfeasibility of $\eta$ at some round $t$ by setting $\mathcal{P}_{t}(\eta)=0$ (i.e., $\eta$ never occurs at round $t$). In the following, we will adopt bold letters to denote the random variables, and non-bold letters to denote their realization (e.g., $\bm{\eta}(v)$ is a random local-state, and $\eta(v)$ denotes a possible realization of $\bm{\eta}(v)$).
Finally, for any round $t\in[T]$, $f_{t}:2^{V}\times H\rightarrow{\mathbb{R}}_{\geq 0}$ is the objective function at round $t$, that associates a non-negative real-value $f_{t}(S_{t},\eta_{t})$ with each subset of items $S_{t}\subseteq V$ and state $\eta_{t}$. Note that objective functions are allowed to be different at each time step. Such a variability in both the state distribution and objective function, immediately makes inefficient to allocate the same budget to each round: indeed, it is immediate to build a setting in which many rounds may lead to very poor outcomes, and only few of them may lead to a very large outcome, and hence all the budget allocated to the former rounds is wasted (see Remark 1 for further details).
Let $\bar{f}:(2^{V}\times H)^{T}\rightarrow{\mathbb{R}}_{\geq 0}$ be the aggregated objective function, that assigns the value $\bar{f}(\vec{S},\vec{\eta}):=\sum_{t=1}^{T}f_{t}(S_{t},\eta_{t})$ to each aggregated subset $\vec{S}:=(S_{1},\ldots,S_{T})$ and aggregated state $\vec{\eta}:=(\eta_{1},\ldots,\eta_{T})$, i.e., it returns the cumulative value of the objective function $f_{t}$ over all rounds $t\in[T]$.
We consider then the problem of selecting an aggregated subset $\vec{S}=(S_{1},S_{2},\ldots,S_{T})$ such that $\bar{f}(\vec{S},\vec{\eta})$ is maximized subject to the budget constraint $\sum_{1}^{T}|S_{t}|\leq B$, for a fixed aggregated state $\vec{\eta}$ drawn from the aggregated state distributions $(\bm{\eta}_{1},\ldots,\bm{\eta}_{T})$. Unfortunately, we need to solve this problem without knowing the aggregated state, and we have access to the state distributions and objective functions only.